Required information E11-4 (Algo) Reporting Stockholders Equity LO11-1, 11-2, 11-3 [The following information applies to theRequired information E11-4 (Algo) Reporting Stockholders Equity LO11-1, 11-2, 11-3 [The following information applies to theRequired information E11-4 (Algo) Reporting Stockholders Equity LO11-1, 11-2, 11-3 [The following information applies to the
Answer #1

Solution 2:

Retained earnings at the beginning of year = Ending retained earnings - Net Income + Dividends

= $800000 - $1130000 + 790000 = $460,000

Solution 3:

Shares in treasury stock = Shares issued - Shares outstanding = 100000 - 76000 = 24,000 shares

Solution 4:

Earnings per share = Net income / Shares outstanding = $1130000 / 76000 = $14.87